# Exploring Mutual Empowerment Between Wireless Networks and RL-based LLMs: A Survey

### v2

@@ IV-C Potential Solutions
Enabling Edge Computing for LLM Training and Inference. Integrating edge computing can enhance the training and inferencing processes in LLMs, resulting in a significant reduction in latency. This approach not only diversifies the dataset by incorporating different data sources but also speeds up response times from LLMs by bringing computation and data storage closer to data sources. Moreover, local data processing on edge devices or devices close to edge devices can minimize the transmission of sensitive information over the internet, thereby enhancing data privacy and security. Additionally, leveraging edge computing can decrease operational expenses by reducing dependence on centralized cloud infrastructure and lowering data transfer costs. This improves the scalability of LLM applications, enabling them to manage larger data volumes and more intricate tasks.
@@
Model Size, Memory, and Computation Reduction. LLMs are usually large and complex, but several techniques can make them smaller and faster without losing performance. Pruning [174] involves cutting out less important parts of the model to reduce its size without much impact on how well it works. Quantization [175] converts high-precision data into lower-precision data, such as changing 32-bit numbers to 8-bit numbers, to save space while maintaining similar performance. Knowledge distillation [176] involves training a smaller model (student) to copy the behavior of a larger model (teacher), effectively transferring knowledge while reducing size. Caching [177] stores intermediate results so they can be reused later, reducing the need for repeated calculations and speeding up the response times of LLMs. By implementing these strategies, LLMs can be integrated into wireless networks while mitigating the complexities associated.
@@
Tool Usage and Prompt Engineering. The challenge of interpretability in LLMs is generating accurate predictions supported by accurate explanations. One approach to address this issue is equipping LLMs with the capability to offer CoT prompt [173], enabling them to break down complex problems into simpler components and reason through intricate scenarios step by step. This CoT can be achieved through the utilization of RL-based techniques, as demonstrated in [28] and discussed in Section III-D. By integrating the CoT capability, LLMs can improve decision-making processes and enhance the reliability of their decisions. This transparency can enhance trust and understanding of the model’s inner workings, increasing acceptance and utilization in various applications. Additionally, the ability of LLMs to break down complex problems and reason through them step by step can significantly benefit users in comprehending the rationale behind the model’s outputs. Beyond CoT, more advanced prompting strategies, such as tree-of-thought (ToT) [178] prompting and graph-of-thought (GoT) [179] prompting, further enhance interpretability. ToT prompting structures reasoning as a tree, enabling LLMs to explore multiple paths, backtrack, and reassess solutions. GoT prompting extends this by using a graph-based approach, allowing flexible, non-linear problem-solving similar to human thinking.
@@
Furthermore, leveraging external tools like the retrieval-augmented generation (RAG) mechanism and external validation sources allows LLMs to verify outputs and enhance information retrieval accuracy. This is expected to reduce hallucination risks while improving response trustworthiness [81]. For example, integrating a mathematical solver can enable precise calculations, while code execution environments allow models to generate and verify programmatic outputs before returning a response. Carefully designed prompts can further reinforce step-by-step reasoning, ensuring that the model’s outputs are not only correct but also explainable.
Leveraging Privacy-Preserving Technologies. Ensuring data privacy and security when applying AI to the industry is indeed important. Training LLMs require vast amounts of data, and storing data centrally poses significant risks, making it easier for hackers to attack and extract user information. To address this issue, federated learning [14, 129] is a promising approach that maintains data privacy by keeping data on individual devices rather than transferring it to a central server. In federated learning, the model is trained locally on each device using the device’s data, and only the model updates are shared with the central server. This prevents sensitive data from leaving the device, thus protecting user privacy. Furthermore, methods that avoid centralized data storage can reduce the risk of single-point attacks by attackers. Technologies such as secure multi-party computation (SMPC) and homomorphic encryption can be integrated with federated learning to enhance privacy and security. Specifically, SMPC allows multiple parties to jointly compute functions of their inputs while maintaining the privacy of those inputs. Meanwhile, homomorphic encryption enables computations to be performed directly on encrypted data without decryption, ensuring the security of the entire data chain.
In addition, split learning [180] is another powerful approach that can bolster data privacy and security. By distributing the training process between edge devices and a central server, split learning reduces the risk of data leakage. In this method, only the intermediate layers of the model are shared with the server, while the raw data remains on the edge devices. This approach significantly minimizes the chances of sensitive information being intercepted or decrypted through simple extraction techniques. An overview of potential solutions is illustrated in Fig. 9.
@@
Figure 10: Motivations, challenges, and potential solutions for empowering LLMs with wireless communication systems.
V Wireless Networks for RL-based LLMs
@@
In this section, we outline the key motivations for leveraging wireless communication systems in reinforcement learning (RL)-based large language models (LLMs), including the limited access to large-scale compliant data, the democratization of LLM development through wireless networks, and the challenge of model staleness amid the rapid expansion of data. We then explore the challenges of integrating wireless communication systems into RL-based LLMs from the perspectives of data, models, environment, and users. Finally, we discuss potential solutions to address the challenges mentioned above. Fig. 10 offers a comprehensive overview of the motivations, challenges, and potential methods for integrating wireless communication system into LLMs.
@@
Figure 11: Challenges for empowering LLMs with wireless communication systems.
V-A Motivations for Empowering RL-based LLMs with Wireless Networks
@@
Limited Access to Large-Scale Compliant Data for LLMs. Given the large number of parameters in LLMs, training them on vast amounts of high-quality data from sources like the Internet, BookCorpus, Wikipedia, and other large-scale datasets can be prohibitively expensive. For instance, training a model like GPT-3, which has 175 billion parameters, requires substantial computational resources and access to massive datasets, often reaching millions of dollars in costs for infrastructure and energy [48]. However, access to such datasets is often restricted to privacy regulations, such as GDPR [181] or HIPAA [182], posing significant challenges, as many publicly available data sources may not meet these legal and ethical requirements. For example, in healthcare applications, patient records are protected by strict privacy laws, making it difficult to train medical LLMs with diverse, real-world data.
@@
Wireless communication systems, with their distributed nature and continuously advancing hardware and software capabilities, offer a promising solution for achieving decentralization and joint data aggregation across distributed sources. By leveraging technologies like continual learning [183], federated learning [184], and secure multi-party computation [185], model training can be distributed across various edge devices. Each edge device can collaborate to update the model based on its private dataset, eliminating the need for a central organizer to access the original data, thus ensuring privacy-preserving large-scale distributed training. Additionally, new edge devices can join the network at any time, seamlessly providing additional resources to enhance the scalability of model training. For example, in autonomous driving, vehicles can collaboratively train LLMs by sharing model updates over wireless networks without exposing sensitive sensor data. This method not only ensures compliance with data privacy regulations but also enhances the model’s adaptability to real-world conditions.
@@
Democratizing LLM Development through Wireless Networks. It is widely recognized that the investment required for AI research and development is substantial, especially for companies developing large models. For example, Sequoia’s research indicates that AI investment will exceed revenue by over \$120 billion in 2024, potentially rising to \$500 billion in 2025 [186]. Microsoft’s CFO has mentioned that the company may need to wait 15 years or more to recover the tens of billions of dollars invested in LLMs [187]. This underscores the enormous cost involved in training and maintaining such models. Currently, the development and deployment of state-of-the-art LLMs are monopolized by a few well-funded tech companies, limiting the participation of other organizations and stifling technological innovation and democratization.
@@
However, it is increasingly feasible for wireless communication systems, particularly in federated or decentralized environments, to democratize access to resources and data. This shift is expected to enable smaller, capital-driven companies or individual entities to contribute to and drive the development of LLMs. Meanwhile, it is worthwhile to highlight that the vision of democratizing LLM development is becoming increasingly achievable. For instance, [188] proposes a method called DiLoCo for training LLMs in a distributed environment. This approach to collaborative model development and maintenance reduces the technical barriers created by resource concentration and promotes innovation. It also minimizes reliance on large-scale centralized infrastructure, allowing LLM development, maintenance, and access to be more broadly distributed and democratized.
@@
Model Staleness in the Face of Rapidly Expanding Data. Model staleness arises when pre-trained models are trained on outdated datasets and fail to keep pace with the rapid expansion of new data. As fresh information continuously emerges, models relying on outdated training data and parameters struggle to adapt to evolving real-world scenarios. This challenge is particularly pronounced in wireless environments, where conditions are highly dynamic and unpredictable. For instance, in autonomous driving, sensor data collected by vehicles changes constantly due to varying road conditions, weather, and traffic patterns [189]. In such cases, pre-trained models that lack real-time adaptation capabilities may become ineffective in responding to new and unforeseen situations, potentially leading to hazardous incidents or accidents.
@@
Wireless communication systems offer a promising solution to this problem through edge computing [190]. By leveraging edge computing techniques such as on-device computing [191] and computation offloading [192], LLMs can be incrementally fine-tuned with new data, enabling continuous adaptation without requiring retraining from scratch. Moreover, reinforcement learning techniques such as RLHF [34] and RLAIF [139], when integrated into wireless networks, allow models to autonomously update and refine their performance over time based on real-world interactions. This ensures that LLMs stay up to date with the latest data trends, enhancing their performance and reliability in dynamic environments.
@@
Low-Latency Interaction and Service Reliability in LLM Deployments. Traditional LLMs are typically deployed in the cloud and hosted on central servers, requiring users to access them via APIs [48]. This setup often introduces delays, particularly when bandwidth is limited or the model is large. In addtion, cloud servers are vulnerable to outages and cyberattacks, as demonstrated by the recent DDoS attack on DeepSeek [193], which temporarily disrupted user access. In such cases, users who rely heavily on centralized servers may experience high latency or service interruptions, significantly degrading the overall user experience.
@@
Energy Efficiency and Sustainable AI Deployment. The training and inference of LLMs demand substantial computational resources, leading to significant energy consumption and an increased carbon footprint. Training these large-scale models often requires running thousands of GPUs or TPUs for extended periods, consuming vast amounts of electricity and necessitating extensive cooling infrastructure to maintain hardware stability. For instance, the latest Grok-3 [194], released in 2025, required a cumulative training time of 200 million hours on 200,000 NVIDIA H100 GPUs. This immense energy demand not only raises concerns about the sustainability of AI development but also poses challenges for deploying such models in resource-constrained environments. In remote or low-resource regions, the limited power supply may make it impractical to train or run LLMs on traditional infrastructure.
Wireless communication systems, particularly when integrated with techniques like federated learning [195], offer a promising solution by leveraging distributed optimization algorithms to reduce communication overhead and energy consumption. For example, the DiLoCo [188] algorithm enables LLM training on a network of devices with poor connectivity, significantly lowering communication frequency and power usage by performing multiple local optimization steps before transmitting updates. Building on this, Photon [196] further enhances collaborative training under weak connectivity and low-bandwidth conditions, introducing the first cost-effective, global-scale LLM pre-training system that achieves twice the convergence speed of previous approaches like DiLoCo. By harnessing wireless communication to facilitate decentralized LLM training, this approach not only enhances energy efficiency but also supports the sustainable deployment of AI, mitigating environmental impact while ensuring broader accessibility.
V-B Challenges
@@
Data Security, Privacy, and Heterogeneity. To date, all of the studies on RL-based LLMs are trained on a centralized server with a massive computing capacity, which demands data collection across a large number of devices and domains to successfully train the model. However, the data are usually generated at the device sites and often exhibit characteristics of heterogeneity, including domain heterogeneity [55], which is a newly summarized concept for this type of variability. These data need to be transmitted to the server through the wired or wireless environment. As a result, the transmission through wireless channels poses a significant risk of eavesdropping and adversarial attacks, which both create severe situations for data transmission for LLMs. In the first case, an eavesdropper can process the passive sniffing by secretly capturing the signal without any interaction with the network or setting up similar configurations with the legitimate receiver to get the data. Then, the signal can be decoded and used for illegitimate purposes. Additionally, the eavesdropper can further interfere with the receiver’s signal reception, causing it to deviate from its original meaning, which can degrade the data quality and destroy the training process of LLMs. Presumably, we can successfully obtain the data without any adversarial attack or privacy issue, but there will still be problems with the heterogeneity distribution of data. The heterogeneous data distribution indicates the unbalance among different types of data, for example, the data in the medical field is extremely scarce, while the data related to sensor networks in smart cities is redundant.
Figure 12: Potential solutions for empowering LLMs with wireless communication systems.
@@
Model Size, Security, and Computational Complexity. As discussed in the motivation section, by leveraging the distribution of devices across physical space, LLMs can be sent down and trained on local devices, thereby reducing the potential risks of transmission of the training data over the wireless network. Nevertheless, the approach consists of several significant drawbacks, such as the large size of the LLMs model (bandwidth consumption), the vulnerability to attacks, and the computing capacity of local devices. In distributed learning, the model is transmitted instead of the data, and this approach usually lowers the communication cost; however, the size of the LLMs is much larger than the normal cases, and thus, it still encounters difficulty when the communication resource is restrained. Another disadvantage of distributing the LLMs to local devices is the malicious training agents, whose target is to degrade the performance of LLMs by attributing their useless weights and updates to the global model aggregation at the central server. Lastly, the dynamic computing capability of local devices should be considered; specifically, this can maximize the utilization of power devices and go easy on the low-computing ones. In real-world scenarios, the approach leads to sustainable and cost-effective deployments by putting more workload and trust in reliable and high-computing agents, while the unimportant tasks should be trained by low-computing agents.

### v3

@@ 3.3.1. Enabling Edge Computing for LLM Training and Inference
Integrating edge computing (Zhang et al., 2024c) can enhance the training and inferencing processes in LLMs, resulting in a significant reduction in latency. This approach not only diversifies the dataset by incorporating different data sources but also speeds up response times from LLMs by bringing computation and data storage closer to data sources. Moreover, local data processing on edge devices or devices close to edge devices can minimize the transmission of sensitive information over the internet, thereby enhancing data privacy and security (Zhang et al., 2018). Additionally, leveraging edge computing can decrease operational expenses by reducing dependence on centralized cloud infrastructure and lowering data transfer costs (Zhao et al., 2019). This improves the scalability of LLM applications, enabling them to manage larger data volumes and more intricate tasks.
3.3.2. Model Size, Memory, and Computation Reduction
@@
LLMs are usually large and complex, but several techniques can make them smaller and faster without losing performance. Pruning (Ashkboos et al., 2024) involves cutting out less important parts of the model to reduce its size without much impact on how well it works. Quantization (Shao et al., 2024a) converts high-precision data into lower-precision data, such as changing 32-bit numbers to 8-bit numbers, to save space while maintaining similar performance. Knowledge distillation (Gu et al., 2024) involves training a smaller model (student) to copy the behavior of a larger model (teacher), effectively transferring knowledge while reducing size. Caching (Bang, 2023) stores intermediate results so they can be reused later, reducing the need for repeated calculations and speeding up the response times of LLMs. By implementing these strategies, LLMs can be integrated into wireless networks while mitigating the complexities associated.
3.3.3. Tool Usage and Prompt Engineering
@@
The challenge of interpretability in LLMs is generating accurate predictions supported by accurate explanations. One approach to address this issue is equipping LLMs with the capability to offer CoT prompt (Wei et al., 2022b), enabling them to break down complex problems into simpler components and reason through intricate scenarios step by step. This CoT can be achieved through the utilization of RL-based techniques, as demonstrated in (Guo et al., 2025) and discussed in Section 2.2.3. By integrating the CoT capability, LLMs can improve decision-making processes and enhance the reliability of their decisions. This transparency can enhance trust and understanding of the model’s inner workings, increasing acceptance and utilization in various applications. Additionally, the ability of LLMs to break down complex problems and reason through them step by step can significantly benefit users in comprehending the rationale behind the model’s outputs. Beyond CoT, more advanced prompting strategies, such as tree-of-thought (ToT) (Yao et al., 2024) prompting and graph-of-thought (GoT) (Besta et al., 2024) prompting, further enhance interpretability. ToT prompting structures reasoning as a tree, enabling LLMs to explore multiple paths, backtrack, and reassess solutions. GoT prompting extends this by using a graph-based approach, allowing flexible, non-linear problem-solving similar to human thinking.
@@
Furthermore, leveraging external tools like the retrieval-augmented generation (RAG) mechanism and external validation sources allows LLMs to verify outputs and enhance information retrieval accuracy. This is expected to reduce hallucination risks while improving response trustworthiness (Shao et al., 2024b). For example, integrating a mathematical solver can enable precise calculations, while code execution environments allow models to generate and verify programmatic outputs before returning a response. Carefully designed prompts can further reinforce step-by-step reasoning, ensuring that the model’s outputs are not only correct but also explainable.
@@
Figure 8. Potential solutions for empowering wireless communication systems with LLMs.
3.3.4. Leveraging Privacy-Preserving Technologies
Ensuring data privacy and security when applying AI to the industry is indeed important. Training LLMs requires vast amounts of data, and storing data centrally poses significant risks, making it easier for hackers to attack and extract user information. To address this issue, federated learning (Nguyen et al., 2024b; Jiang et al., 2024) is a promising approach that maintains data privacy by keeping data on individual devices rather than transferring it to a central server. In federated learning, the model is trained locally on each device using the device’s data, and only the model updates are shared with the central server. This prevents sensitive data from leaving the device, thus protecting user privacy. Furthermore, methods that avoid centralized data storage can reduce the risk of single-point attacks by attackers. Technologies such as secure multi-party computation (SMPC) (Goldreich, 1998) and homomorphic encryption (Acar et al., 2018) can be integrated with federated learning to enhance privacy and security. Specifically, SMPC allows multiple parties to jointly compute functions of their inputs while maintaining the privacy of those inputs. Meanwhile, homomorphic encryption enables computations to be performed directly on encrypted data without decryption, ensuring the security of the entire data chain.
In addition, split learning (Thapa et al., 2022) is another powerful approach that can bolster data privacy and security. By distributing the training process between edge devices and a central server, split learning reduces the risk of data leakage. In this method, only the intermediate layers of the model are shared with the server, while the raw data remains on the edge devices. This approach significantly minimizes the chances of sensitive information being intercepted or decrypted through simple extraction techniques. An overview of potential solutions is illustrated in Fig. 8.
4. Wireless Networks for RL-based LLMs
@@
In this section, we outline the key motivations for leveraging wireless communication systems in reinforcement learning (RL)-based large language models (LLMs), including the limited access to large-scale compliant data, the democratization of LLM development through wireless networks, and the challenge of model staleness amid the rapid expansion of data. We then explore the challenges of integrating wireless communication systems into RL-based LLMs from the perspectives of data, models, environment, and users. Finally, we discuss potential solutions to address the challenges mentioned above. Fig. 9 offers a comprehensive overview of the motivations, challenges, and potential methods for integrating wireless communication systems into LLMs.
@@
Figure 9. Motivations, challenges, and potential solutions for empowering LLMs with wireless communication systems.
4.1. Motivations for Empowering RL-based LLMs with Wireless Networks
4.1.1. Limited Access to Large-Scale Compliant Data for LLMs
@@
Given a large number of parameters in LLMs, training them on vast amounts of high-quality data from sources like the Internet, book corpus, Wikipedia, and other large-scale datasets can be prohibitively expensive (Zhao et al., 2024a; Zhang et al., 2024b). For instance, training a model like GPT-3, which has 175 billion parameters, requires substantial computational resources and access to massive datasets, often reaching millions of dollars in costs for infrastructure and energy (Wallstreetcn, 2023). However, access to such datasets is often restricted to privacy regulations, such as GDPR (Voigt and Von dem Bussche, 2017) or HIPAA (Cohen and Mello, 2018), posing significant challenges, as many publicly available data sources may not meet these legal and ethical requirements. For example, in healthcare applications, patient records are protected by strict privacy laws, making it difficult to train medical LLMs with diverse, real-world data.
@@
Wireless communication systems, with their distributed nature and continuously advancing hardware and software capabilities, offer a promising solution for achieving decentralization and joint data aggregation across distributed sources. By leveraging technologies like continual learning (Wang et al., 2024d), federated learning (Qiao et al., 2023f), and secure multi-party computation (Zhou et al., 2024b), model training can be distributed across various edge devices. Each edge device can collaborate to update the model based on its private dataset, eliminating the need for a central organizer to access the original data, thus ensuring privacy-preserving large-scale distributed training. Additionally, new edge devices can join the network at any time, seamlessly providing additional resources to enhance the scalability of model training. For example, in autonomous driving, vehicles can collaboratively train LLMs by sharing model updates over wireless networks without exposing sensitive sensor data. This method not only ensures compliance with data privacy regulations but also enhances the model’s adaptability to real-world conditions.
4.1.2. Democratizing LLM Development through Wireless Networks
@@
It is widely recognized that the investment required for AI research and development is substantial, especially for companies developing large models (Mou, 2019). For example, Sequoia’s research indicates that AI investment will exceed revenue by over \$120 billion in 2024, potentially rising to \$500 billion in 2025 (Widder and Hicks, 2024). Microsoft’s CFO has mentioned that the company may need to wait 15 years or more to recover the tens of billions of dollars invested in LLMs (Cohan, 2024). This underscores the enormous cost involved in training and maintaining such models. Currently, the development and deployment of state-of-the-art LLMs are monopolized by a few well-funded tech companies, limiting the participation of other organizations and stifling technological innovation and democratization.
@@
However, it is increasingly feasible for wireless communication systems, particularly in federated or decentralized environments, to democratize access to resources and data. This shift is expected to enable smaller, capital-driven companies or individual entities to contribute to and drive the development of LLMs. Meanwhile, it is worthwhile to highlight that the vision of democratizing LLM development is becoming increasingly achievable. For instance, (Douillard et al., 2023) proposes a method called DiLoCo for training LLMs in a distributed environment. This approach to collaborative model development and maintenance reduces the technical barriers created by resource concentration and promotes innovation. It also minimizes reliance on large-scale centralized infrastructure, allowing LLM development, maintenance, and access to be more broadly distributed and democratized.
4.1.3. Model Staleness in the Face of Rapidly Expanding Data
@@
Model staleness arises when pre-trained models are trained on outdated datasets and fail to keep pace with the rapid expansion of new data (Chen et al., 2022). As fresh information continuously emerges, models relying on outdated training data and parameters struggle to adapt to evolving real-world scenarios. This challenge is particularly pronounced in wireless environments, where conditions are highly dynamic and unpredictable. For instance, in autonomous driving, sensor data collected by vehicles changes constantly due to varying road conditions, weather, and traffic patterns (Zhang et al., 2023). In such cases, pre-trained models that lack real-time adaptation capabilities may become ineffective in responding to new and unforeseen situations, potentially leading to hazardous incidents or accidents.
@@
Wireless communication systems offer a promising solution to this problem through edge computing (Qiao et al., 2023e). By leveraging edge computing techniques such as on-device computing (Pandey et al., 2020) and computation offloading (Hossain et al., 2023), LLMs can be incrementally fine-tuned with new data, enabling continuous adaptation without requiring retraining from scratch. Moreover, reinforcement learning techniques such as RLHF (Ouyang et al., 2022) and RLAIF (Bai et al., 2022), when integrated into wireless networks, allow models to autonomously update and refine their performance over time based on real-world interactions. This ensures that LLMs stay up to date with the latest data trends, enhancing their performance and reliability in dynamic environments.
4.1.4. Low-Latency Interaction and Service Reliability in LLM Deployments
@@
Traditional LLMs are typically deployed in the cloud and hosted on central servers, requiring users to access them via APIs (Wallstreetcn, 2023). This setup often introduces delays, particularly when bandwidth is limited or the model is large. In addition, cloud servers are vulnerable to outages and cyberattacks, as demonstrated by the recent DDoS attack on DeepSeek (Xu et al., 2025a), which temporarily disrupted user access. In such cases, users who rely heavily on centralized servers may experience high latency or service interruptions, significantly degrading the overall user experience.
@@
4.1.5. Energy Efficiency and Sustainable AI Deployment
The training and inference of LLMs demand substantial computational resources, leading to significant energy consumption and an increased carbon footprint (Iftikhar and Davy, 2024). Training these large-scale models often requires running thousands of GPUs or TPUs for extended periods, consuming vast amounts of electricity and necessitating extensive cooling infrastructure to maintain hardware stability. For instance, the latest Grok-3 (Reyes-Rivera et al., 2024), released in 2025, required a cumulative training time of 200 million hours on 200,000 NVIDIA H100 GPUs. This immense energy demand not only raises concerns about the sustainability of AI development but also poses challenges for deploying such models in resource-constrained environments. In remote or low-resource regions, the limited power supply may make it impractical to train or run LLMs on traditional infrastructure.
Wireless communication systems, particularly when integrated with techniques like federated learning (Qiao et al., 2023a), offer a promising solution by leveraging distributed optimization algorithms to reduce communication overhead and energy consumption. For example, the DiLoCo (Douillard et al., 2023) algorithm enables LLM training on a network of devices with poor connectivity, significantly lowering communication frequency and power usage by performing multiple local optimization steps before transmitting updates. Building on this, Photon (Sani et al., 2024) further enhances collaborative training under weak connectivity and low-bandwidth conditions, introducing the first cost-effective, global-scale LLM pre-training system that achieves twice the convergence speed of previous approaches like DiLoCo. By harnessing wireless communication to facilitate decentralized LLM training, this approach not only enhances energy efficiency but also supports the sustainable deployment of AI, mitigating environmental impact while ensuring broader accessibility.
4.2. Challenges
Figure 10. Challenges for empowering LLMs with wireless communication systems.
@@ 4.2.1. Data Security, Privacy, and Heterogeneity
To date, all of the studies on RL-based LLMs are trained on a centralized server with a massive computing capacity, which demands data collection across a large number of devices and domains to train the model successfully (Ren et al., 2024). However, the data are usually generated at the device sites and often exhibit characteristics of heterogeneity, including domain heterogeneity (Qiao et al., 2025b), which is a newly summarized concept for this type of variability. These data need to be transmitted to the server through the wired or wireless environment. As a result, the transmission through wireless channels poses a significant risk of eavesdropping and adversarial attacks (Qiao et al., 2024d), which both create severe situations for data transmission for LLMs. In the first case, an eavesdropper can process the passive sniffing by secretly capturing the signal without any interaction with the network or setting up similar configurations with the legitimate receiver to get the data. Then, the signal can be decoded and used for illegitimate purposes. Additionally, the eavesdropper can further interfere with the receiver’s signal reception, causing it to deviate from its original meaning, which can degrade the data quality and destroy the training process of LLMs. Presumably, we can successfully obtain the data without any adversarial attack or privacy issue, but there will still be problems with the heterogeneity distribution of data. The heterogeneous data distribution indicates the unbalance among different types of data. For example, the data in the medical field is extremely scarce, while the data related to sensor networks in smart cities is redundant (Liu et al., 2024c).
4.2.2. Model Size, Security, and Computational Complexity.
@@
As discussed in the motivation section, by leveraging the distribution of devices across physical space, LLMs can be sent down and trained on local devices, thereby reducing the potential risks of transmission of the training data over the wireless network. Nevertheless, the approach consists of several significant drawbacks, such as the large size of the LLMs model (bandwidth consumption), the vulnerability to attacks, and the computing capacity of local devices (Hasan et al., 2024). In distributed learning, the model is transmitted instead of the data, and this approach usually lowers the communication cost; however, the size of the LLMs is much larger than the normal cases, and thus, it still encounters difficulty when the communication resource is restrained (Ren et al., 2024). Another disadvantage of distributing the LLMs to local devices is the malicious training agents, whose target is to degrade the performance of LLMs by attributing their useless weights and updates to the global model aggregation at the central server (Chen et al., 2022). Lastly, the dynamic computing capability of local devices should be considered; specifically, this can maximize the utilization of power devices and go easy on the low-computing ones. In real-world scenarios, the approach leads to sustainable and cost-effective deployments by putting more workload and trust in reliable and high-computing agents, while the unimportant tasks should be trained by low-computing agents.
4.2.3. Stability and Reliability of Network Environments